Check your understanding Dan is calculating the opportunity cost of getting his college degree. If he goes to college, each year, he’ll pay $9,000 for tuition, $1,400 for textbooks, $4,500 for housing, and $2,500 for food. If he decides not to go to college, his parents will let him live at home for free, and he will spend only $1,000 per year on food. The rest of his expenses won’t depend on whether or not he goes to college. While deciding whether to go to college, Dan is offered a job paying $30,000 per year, which he won’t be able to take if he goes to college. What is Dan’s explicit cost of a year in college ($)? What is Dan’s opportunity cost of four years in college ($)?
Answer #1

Answer

Explicit cost is the direct payment made for the alternative you have chosen.

Here He chose to go to school. Hence Explicit cost is the cost he incur to spend 1 year at school i.e. $9,000 for tuition, $1,400 for textbooks, $4,500 for housing, and $2,500 for food. Hence Total explicit cost of Dan for a year in college is given by:

Explicit Cost = 9000 + 1400 + 4500 + 2500 = 17400

Hence, Dan’s explicit cost of a year in college = $17400

Dan opportunity cost of 4 years in college is the sum of implicit cost and explicit cost for 4 year.

Dan’s explicit cost of 4 years in college = $17400*4 = $69,600

Implicit cost is the net income he sacrifice if he didn't went for college.

That Net income his Total income from Job for 4 years - his total expenditure on food for fous years at home.

Hence Implicit cost = 30,000*4 - 1000*4 = 116000

Hence Opportunity cost of four years in college = Total Implicit cost + Total explicit cost

= 116000 + 69,600 = 185600

Hence,  Dan’s opportunity cost of four years in college = $185,600